OpenStudy (anonymous):

The cube has a mass of 3300 g. What is its density? g/cm3 What substance is it? 30cm=length 10cm=width 50cm=height

OpenStudy (anonymous):

30*10*50= 300*50=15000 3300/15000 33/150 11/50 0.22

OpenStudy (saifoo.khan):

0.22g/cm^3

OpenStudy (anonymous):

V= 30x10x50 = 15000 cm^3 D= m/v= 3300/15000 = 0.22 g/cm^3 substance, i'm not so sure, get to look it up a lil bit online